OpenStudy (anonymous):

in a geometric progression,the sum of the second and third term is 9,and the seventh term is eight times the fourth.find the first term,the common ratio and the fifth term

OpenStudy (anonymous):

u7=8u4 ar^6=8ar^3 r=2 u2+u3=9 ar(1+r)=9 2a(3)=9 a=3/2 u5=ar^4=3/2x16=24